Both GSI vacuum group and edinburgh ion gauges OFF
Fees power cycled. All system checks OK.
Writing to file, alpha background overnight.
Moved detector IN 2mm. Noise appears reduced with any distance moved in so overnight spectra to be taken while IN. Not moving further in as possible beam tomorrow.
22:44
Leakage current stable, slowly decreasing ~3.57 uA (attachment 1)
Statistics look OK
Temperatures OK
System wide check all pass - OK
Noise appears much reduced on spectrum Browser channels. (attachement 2). Compared to same plot on elog 197
XYtotal per pixel (attachment 3)
Ex histogram (attachment 4)
|